OpenStudy (anonymous):

how do I expand this equation y=x^​2-6x+8

OpenStudy (anonymous):

(x-2)(x-4)

OpenStudy (karatechopper):

If we factor it out. then you would start by making two sets of parenthesis. y= ( ) ( ) So, to start out, lets put x^2 into the parenthesis. For better understanding, we know that, x^2= x multiplied by x so we can put an x in both of the parenthesis. y= ( x ) ( x ) Great! Now that we have taken care of the x's we can now focus on the two factors that multiply to equal 8. And add to become -6. Lets start that by listing the factors of 8: 1,8; 2,4; -1,-8 ; -2,-4 We only have 4 pairs that multiply to become 8. Only one pair of those 4 can add to become -6. Lets see which one. 1+8=9 2+4=6 -1+ (-8)= -9 -2 + (-4)= -6 Looks like only the -2 and -4 will work. Lets insert that into the parenthesis now! y= ( x -2 ) ( x-4 ) Now we have factored! If that was what you meant by "expanding"

OpenStudy (anonymous):

yh same thing ha! ??

OpenStudy (karatechopper):

But I showed my work, didn't give the answer.